# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v4

@@ 6 Discussion
The effectiveness of our framework remains dependent on the quality and diversity of accessible data, and its performance could differ in alternate contexts. Additionally, while the incorporation of explainability techniques contributes to interpretability, it might not encompass all underlying intricacies of network behavior. The selection of the two local methods necessitates further exploration of alternative approaches to ensure comprehensiveness. The enhancement of classification outcomes demands the incorporation of additional information beyond structural MRI data to achieve improved accuracy. To further investigate and incorporate more intricate strategies, the combination of XAI techniques and reduce dimensionality methods can explore overlapping aspects. Furthermore, complex approaches like t-SNE for non-linear dimensionality reduction can be employed to delve deeper into the subject.
Moving forward, we envision expanding the framework’s applicability to other neurological conditions and classification tasks (such as schizophrenia or bipolar disorder), employing other external datasets and further refining the interpretability techniques. Our study paves the way for a more systematic exploration of sulcal variability through the lens of deep learning, which can help explore cognitive and functional variability as well as pathological changes.
7 Conclusion
In this study, we classified the presence and absence of the secondary sulcus, the PCS, in a highly heterogeneous, unbiased, and well-annotated cohort of 596 structural MRIs. We used traditional 3D deep convolutional neural and attention mechanism networks to achieve acceptable detection accuracy and apply explainability techniques for new-pattern discovery.
As there were no global explainability methodologies in the literature efficient enough to deliver robust, accurate, low-complexity and three-dimensional explanations, we developed and proposed a novel XAI 3D-Framework to provide global explanations for the task of PCS detection. The global explanations derived from this framework outperformed established methods such as SHAP and GradCam in terms of faithfulness, achieving the quality needed to discover new global patterns. Additionally, the 3D representation of the proposed framework allows for the robust identification of complex new patterns within the high complexity and dimensionality of neuroscience applications. To accomplish this, the 3D-Framework enables the identification of sub-regions that are critical to decision-making through a post-fusion transformation of global explanations and statistical features.
The detection of the presence of the PCS appeared to focus most reliably on features related to the posterior temporal region and the internal parietal region (on the sulcal skeleton), with additional focus on the cingulate region and thalamus when focusing on the grey-white surface. This highlights potential co-variation between these structures which should be investigated further, notably in terms of common genetic influences in these regions, both for neurodevelopmental and functional or pathological considerations.
Significantly, our framework has revealed notable correlations between specific brain sub-regions and the presence or absence of the paracingulate sulcus, offering deeper insights into underlying developmental mechanisms. This approach opens the door to more robust and reliable explanations in medical imaging and other complex computer vision tasks.
This work advances the fields of both deep learning and neuroscience, enabling automated and impartial annotations while illuminating detailed neuroanatomical relationships across diverse subregions. Our XAI 3D-Framework offers unparalleled insights, thus setting the stage for comprehensive exploration within the ever-evolving domains of deep learning and neuroscience.
Acknowledgment

### v5

@@ 4 Discussion
The effectiveness of extracting generalized patterns using our proposed framework underscores the importance of incorporating data from multiple cohorts. To this end, we plan to apply the framework to additional cohorts, such as BeneMin (Deakin et al. (2018)) and Biobank (Sudlow et al. (2015)), to identify shared patterns in the classification of PCS presence and absence. Combining XAI techniques with dimensionality reduction methods may further reveal overlapping aspects of the data. Advanced approaches, such as t-SNE for non-linear dimensionality reduction, could also provide deeper insights into these relationships. Additionally, we aim to extend the framework’s application to other neurological conditions and classification tasks, including schizophrenia and bipolar disorder, by leveraging external datasets and improving interpretability techniques.
This study establishes a foundation for systematic exploration of sulcal variability through deep learning, with the potential to advance our understanding of cognitive and functional variability as well as pathological changes.
4.1 Conclusion
In this study, we classified the presence or absence of the paracingulate sulcus (PCS) in a diverse cohort of 596 structural MRIs using 3D deep convolutional neural networks and attention mechanisms. To address the lack of robust global explainability methods for 3D neuroimaging data, we developed an innovative XAI 3D-Framework. This framework provides accurate and low-complexity global explanations for PCS detection by integrating statistical features (Shape) with XAI methods (GradCam and SHAP) alongside reduced dimensionality information, ensuring that the explanations capture both model learning and cohort-specific variability. Furthermore, the combined application of GradCam and SHAP mitigates inter-method variability, thereby enhancing the reliability and robustness of the explanations.
Our framework outperformed established methods like GradCam and SHAP in faithfulness, enabling the robust identification of sub-regions critical for decision-making through a fusion of global explanations and statistical features. Key patterns identified include a focus on the posterior temporal and internal parietal regions on the sulcal skeleton, and on the cingulate region and thalamus when analyzing the grey-white surface. These findings indicate potential co-variation between these structures, likely underpinned by shared genetic or developmental mechanisms. Such insights hold significant implications for both neurodevelopmental and pathological research, providing a foundational framework for guiding future investigative trajectories.
Our work advances both deep learning and neuroscience by enabling automated, unbiased annotations and delivering unprecedented insights into sulcal variability and its functional or pathological relevance. The XAI 3D-Framework sets the stage for broader applications in medical imaging and other complex computer vision tasks, providing a foundation for comprehensive exploration of neuroanatomy and developmental mechanisms.
Data and Code Availability
Data availability
Code availability
Author Contributions
M.M, H.V., P.L., J.S. and G.M. conceived the study. M.M. wrote the code and conducted the experiments. I.A., LE.MJ., H.V., A.AM., and S.C.M. collected and pre-processed the data cohort. M.M, and H.V. analyzed the data and results. M.M. contributed to pulling deep learning and XAI methods and conducted chart reviews. M.M contributed to the experimental design and validation protocol. G.M.,J.S.,P.L., H.V. and M.M. was in charge of overall direction and planning. All authors contributed to the interpretation of the results. M.M., H.V. and P.L. visualized the study and extracted figures, and M.M and H.V. drafted the manuscript, which was reviewed, revised and approved by all authors.
Funding
This study was supported by funding from the Medical Research Council, grant number: MR/W020025/1.
Declaration of Competing Interests
GKM consults for ieso digital health. All other authors declare that they have no competing interests.
Acknowledgements